a)
) In a certain market, the demand for peach was given as QD = 400 -3P, the
first day of marketing in 2020, in August 2020, the demand for peach is now
given as QD = 200 -3P. From the statement above tell in one sentence the
change in demand; what did you consider as the indicator? Support your answers by
sketching the equations above on the same graph sheet.
b) In a certain market for Sobolo when the price of the drink was GHC 3 per
bottle, the quantity demanded was 12 bottles. On another day, when the price
per bottle was GHC 5, quantity demanded was 6. Use the information to write
down and equation for demand. Now, what will be quantity demanded if price
per bottle was 6? If we restrict price to the same demand curve, what will
happen if price is less than GHC 2?
c) Complementary in supply are goods for which producing more of one
requires producing more of the other. Using the definition, write down any 5
real life examples of complements in supply
1
Expert's answer
2021-08-17T10:23:38-0400

(a) "Q_D=400-3P"

"Q_D=200-3P"

The change in demand is contractionary.

This is indicated by shift of the demand curve to the left.

(b)

"P_1=3, Q_1=12"

"P_2=5,Q_2=6"

The demand equation is given by the form:

"Q=a+mP"

"m=slope =\frac {Q_2-Q_1}{P_2-P_1}"

"=\frac {6-12}{5-3}=-3"

"Q=a-3P"

substituting Q=12 and P=3

we get

"a=12+9=21"

The demand equation will be:

"Q=21-3P"


If P=6, the quantity demanded will be:

"Q=21-3\times6=3 bottles"


If the price is less than 2 GHC the quantity demanded will rise to a higher level.

"Q=21-3P"

"\implies""P=7-\frac {Q}{3}"

given that P <2,

"\implies" "7-\frac{Q}{3}<2"

"15<Q"

If the price is less than 2, then the quantity demanded will be greater than 15.

(c) Real life examples of compliments in supply

  • Petrol and cars
  • Pencils and notebooks
  • Mobile phones and sim cards
  • Tennis balls and tennis racket
  • Burger and Burger Buns.